OpenStudy (anonymous):

Determine whether a triangle can be formed with the given side lengths. If so, use Heron's formula to find the area of the triangle. a = 240 b = 127 c = 281

OpenStudy (anonymous):

check if sum of any two side is always greater than the reamining side...

OpenStudy (anonymous):

Heron's formula for the area of triangle is given by sqrt( s*(s-a)*(s-b)*(s-c)) where 2s=a+b+c

OpenStudy (anonymous):

Heron say it look : p=1/2 all the p 2p= all the p Heron says : \[\sqrt{p(p-a)(p-b)(p-c)}\]

OpenStudy (anonymous):

I did that, but I'm not getting anything equivalent to my choices.

OpenStudy (anonymous):

u didnt give numbers !

OpenStudy (anonymous):

@Loveiskey18

OpenStudy (anonymous):

@@@@@@@Loveiskey18

OpenStudy (anonymous):

My choices: No triangle is formed. 21,237.83 21,223.73 15,183.77

OpenStudy (anonymous):

My apologies.

OpenStudy (anonymous):

My nean was a , b ,c in triangle ! u didnt give them!

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

a = 240 b = 127 c = 281

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

Ok ! Now find p : p=240+127+281/2 OK ?!

OpenStudy (anonymous):

324.

OpenStudy (anonymous):

Exelent ! Now go with formula : (get a , b , c , p)in and get numbers ! Can u do it ?!:)

OpenStudy (anonymous):

What is P?

OpenStudy (anonymous):

Oh excuse me its 2p means around !

OpenStudy (anonymous):

2p maens that : 240+127+281/2 . I was falled last :( Now can u answer ? Look : in heron we want p means 1/2 2p . OK ?!

OpenStudy (anonymous):

So for p, I put 2?

OpenStudy (anonymous):

Yeah ! So p=324/2=162

OpenStudy (anonymous):

@Loveiskey18 :Can u answer friend ?

OpenStudy (anonymous):

I do not know what to do after You did that:p=324/2=162

OpenStudy (anonymous):

OK ! Now we have the Heron formula : \[\sqrt{p(p-a)(p-b)(p-c)}\]

OpenStudy (wolf1728):

Another way to express the formula where s is semi-perimeter Semi-Perimeter = (240+127+281)/2 = 648/2 = 324 area = square root (s • (s - 240) • (s - 127) • (s - 281))

OpenStudy (anonymous):

Si its : \[\sqrt{162(162-240)(p-127)(162-281)}\]

OpenStudy (anonymous):

@wolf1728:u are realy fall friend ! It s p not s .!

OpenStudy (anonymous):

@Loveiskey18 : Got it ?!:)

OpenStudy (anonymous):

Yes. I got 2923830 sqrt(2)

OpenStudy (anonymous):

Any problem ? Did u get it rely ?!:)

OpenStudy (wolf1728):

I calculate 15,183.765540866336 or 15,184 rounded

OpenStudy (anonymous):

The answer isnt importent I dont calculat that ! Formula is important !:)

OpenStudy (anonymous):

Yes, I obtained the answer using Heron's formula

OpenStudy (anonymous):

OK :)

OpenStudy (anonymous):

@E.ali Why do you have a smiley face, according to @wolf1728 my answer is incorrect.

OpenStudy (anonymous):

@Loveiskey18 : I never say that :) I have another bean :(

OpenStudy (anonymous):

What is your bean?

OpenStudy (anonymous):

Thank You both, @wolf1728 and @E.ali

OpenStudy (anonymous):

Your welom Sir !:)

OpenStudy (wolf1728):

Okay and thank you Loveiskey18

OpenStudy (anonymous):

I'm not a Sir.

OpenStudy (anonymous):

Oh excuse me Modom :( I didnt know :( I m sorry

OpenStudy (wolf1728):

Not to run this into the ground but try this calculator http://www.mathsisfun.com/geometry/herons-formula.html or this one http://keisan.casio.com/exec/system/1223267646 or this one http://mste.illinois.edu/dildine/heron/triarea.html and see what you get

OpenStudy (anonymous):

@wolf1728 : Thanks !

OpenStudy (anonymous):

They are all correct, 15183.7.

OpenStudy (anonymous):

:)

OpenStudy (wolf1728):

Thanks Loveiskey18 :-) and Eali :-)

OpenStudy (anonymous):

You're Welcome.

OpenStudy (anonymous):

You re welcome !:)